### v1

# draft_biblio.bib

Abstract.
We consider the least-deletion map on the Cantor space, namely the map that changes the first $1$ in a binary sequence to $0,$ and construct product measures on $2^{\mathbb{N}}$ so that the corresponding Radon–Nikodym cocycles oscillate or converge to zero nonsummably along the forward geodesic of the map. These examples answer two questions of Tserunyan and Tucker-Drob. We analyze the oscillating example in terms of random walks on $\mathbb{Z}$, using the Chung–Fuchs theorem.
Key words and phrases:
Radon–Nikodym cocycle, countable-to-one Borel function, countable Borel equivalence relations, quasi-pmp, measure class preserving, non-singular, random walk, forward geodesic, forward orbit, oscillation
2020 Mathematics Subject Classification:
Primary 03E15, 37A40; Secondary 60G50, 37A20
S. Bell was partially supported by an NSERC Undergraduate Student Research Award, with supplemental funding from FRQNT. O. Rodgers and T. Chu were partially supported by McGill Science Undergraduate Research Awards. The authors were partially supported by Anush Tserunyan’s NSERC Discovery Grant RGPIN-2020-07120.
1. Introduction
This paper focuses on descriptive dynamics of countable Borel equivalence relations (CBERs) on a standard Borel space $X$. Of particular interest are those equivalence relations which admit an acyclic graphing, called a treeing. Such CBERs are called treeable. In the probability-measure-preserving (pmp) setting, Adams’ dichotomy gives a complete characterization of which treeings are amenable: a pmp treeing is amenable if and only almost every connected component has at most two ends.
This dichotomy fails in general in the measure-class-preserving (mcp) setting, but Tserunyan and Tucker-Drob proved the following generalized version of the Adams dichotomy. Their characterization of amenable treeings is in terms of the Radon–Nikodym cocycle associated to the equivalence relation and the measure: if $E$ is an mcp CBER on a standard probability space $(X,\mu)$, then $E$ admits a unique Borel (up to null sets) Radon–Nikodym cocycle $\mathfrak{w}:E\to\mathbb{R}^{+}$ which quantifies the failure of invariance of the measure $\mu$. We think of $\mathfrak{w}_{y}(x):=\mathfrak{w}(x,y)$ as the weight of $x$ relative to the weight of $y$.
Tserunyan and Tucker-Drob characterize the amenability of an mcp treeing via the behaviour of the corresponding Radon–Nikodym cocycle. Namely, they show in [TTD] that a locally countable acyclic mcp graph $T$ is amenable if and only if almost every connected component $C$ has at most two nonvanishing ends. Here, an end $\eta\in\partial_{T}{C}$ is called nonvanishing if for some $x\in C$, there is $\varepsilon>0$ such that for every neighbourhood of the end, there is a point $y$ in that neighbourhood such that $\mathfrak{w}_{x}(y)\geq\varepsilon$. Their analysis of the behaviour of the Radon-Nikodym cocycle along the ends relies on a close examination of the case when the equivalence relation is given by a function.
In this context, they ask the following two questions: is there an example of a measure space $(X,\mu)$ equipped with a Borel transformation $f:X\to X$ such that the induced Radon–Nikodym cocycle oscillates along the $f$-forward geodesic, meaning it gets arbitrarily large and arbitrarily close to zero? Is there an example where the cocycle goes to zero along the forward geodesic, but the infinite sum of the cocycle weights is infinite? We answer both questions in the affirmative.
Acknowledgements
We are grateful to Anush Tserunyan for her mentorship and helpful comments. We especially thank Louigi Addario-Berry for sketching the probabilistic proof that one of the cocycles oscillates as required. We thank Anush Tserunyan and Robin Tucker-Drob for suggesting the least-deletion map as a candidate to analyze.
2. Preliminaries
Let $(X,\mu)$ be a standard probability space and let $Y$ be a standard Borel space. Given a $\mu$-measurable function $f:X\rightarrow Y$, we denote the pushforward of $\mu$ under $f$ by $f_{*}\mu$ where $f_{*}\mu(A)=\mu(f^{-1}(A))$ for each Borel set $A\subseteq Y.$ For a fixed element $x\in X$, we call the path $(x,fx,f^{2}x,\cdots)$ the forward geodesic of $x$.
We say $g:X\rightarrow X$ is measure-class-preserving (or non-singular) if $g_{*}\mu$ is equivalent to $\mu$, i.e. they are both absolutely continuous with respect to each other.
Let $E$ be an equivalence relation. We say $E$ is a countable Borel equivalence relation if each equivalence class $[x]_{E}$ is countable and if the subset $E\subseteq X\times X$ is a Borel subset of $X\times X$. In this case, notice $E$ itself is a standard Borel space.
For a subset $A$ of a standard space $X$ endowed with a countable Borel equivalence relation $E$, we denote the saturation of $A$ by $[A]_{E}=\{x\in X:xEa\text{ for some }a\in A\}.$
We say a measure $\mu$ defined on the Borel sets of $X$ is $E$-quasi-invariant if $[A]_{E}$ is $\mu$-null whenever $A$ is $\mu$-null. Equivalently in this case, we say that the measured equivalence relation $E$ on $(X,\mu)$ is measure-class-preserving (mcp).
By [kechrismiller, Section 8], if $E$ is a CBER on a standard probability space $(X,\mu)$ and $\mu$ is a $E$-quasi-invariant measure, then $E$ admits a unique Borel (up to null sets) Radon–Nikodym cocycle $\mathfrak{w}:E\to\mathbb{R}^{+}$ satisfies the (tilted) mass transport principle, meaning for any Borel partial bijection $\gamma:X\rightharpoonup X$ and any $f\in L^{1}(X,\mu),$
[…]
where $\mathfrak{w}_{x}(\gamma(x))\coloneqq\mathfrak{w}(x,\gamma(x))$.
A more comprehensive reference for the dynamics of countable Borel equivalence relations can be found in [kechrismiller].
3. The least-deletion map and the corresponding Radon–Nikodym cocycle
Our setting is that of a countable Borel equivalence relation $E$ on the Cantor space $2^{\mathbb{N}}$ equipped with an $E$-quasi-invariant probability measure $\mu$.
Let $X$ be the (co-countable) subset of $2^{\mathbb{N}}$ of sequences containing infinitely many $1$s. Consider now the least-deletion map $f:X\rightarrow X$ which flips the first $1$ in a sequence $x=(x_{0},x_{1},\cdots)\in 2^{\mathbb{N}}$ to a 00. Note that $f$ is finite-to-one.
We denote by $E_{f}$ the orbit equivalence relation of $f$; note that $E_{f}$ is generated by the involutions $b_{n}:X\rightarrow X$ which flip the $n$th bit. This implies that any product measure $\mu=\prod_{n=1}^{\infty}m_{n}$ on $2^{\mathbb{N}}$ with non-trivial marginal measures $m_{n}$ is $E_{f}$-quasi-invariant since each of these involutions preserves the measure class of $\mu$ by definition. For any such $\mu,$ we compute the Radon–Nikodym cocycle $\mathfrak{w}:E_{f}\rightarrow\mathbb{R}^{+}$ explicitly, and we derive a formula for $\mathfrak{w}_{x}(f^{k}(x))$ for $k\in\mathbb{N}$.
Proposition 3.1.
Fix a non-trivial product measure $\mu=\prod_{n=1}^{\infty}m_{n}$ on $2^{\mathbb{N}}$, and let $\mathfrak{w}$ be the Radon–Nikodym cocycle associated to $E_{f}$ and $\mu$. Then for each $k\geq 1$ and for a.e. $x\in X$, we have
$\mathfrak{w}_{x}(f^{k}(x))=\prod_{i=0}^{k-1}\dfrac{m_{n_{i}}(0)}{m_{n_{i}}(1)},$
where $(n_{i})_{i\geq 0}$ is the subsequence of indices $n_{i}$ for which $x_{n_{i}}=1$.
Proof.
It is enough to show that for a bit-flip involution $b_{n}$, we have that
$\mathfrak{w}_{x}(b_{n}(x))=\frac{1-m_{n}(x_{n})}{m_{n}(x_{n})}.$
Then, using the cocycle property and that $f$ always flips a one, we get the desired formula. It suffices to show this formula holds on cylinder sets.
Consider a cylinder set $C_{n}=[a_{0}\cdots a_{n}]$ with bits $a_{i}\in\left\{0,1\right\}$ specified up to the nth position. We have $\mu(C_{n})=m_{0}(a_{0})\cdot\ldots\cdot m_{n}(a_{n})$. Then $b_{n}(C_{n})=[a_{0}\ldots(1-a_{n})]$ so $\mu(b_{n}(C_{n}))=m_{0}(a_{0})\cdot\ldots\cdot m_{n-1}(a_{n})\cdot(1-m_{n}(a_{n}))$. Computing the Radon–Nikodym derivative of $b_{n}$ on this cylinder set, we obtain
$\frac{d(b_{n\ast}\mu)}{d\mu}(C_{n})=\frac{m_{0}(a_{0})\cdot\ldots\cdot(1-m_{n}(a_{n}))}{m_{0}\cdot\cdot\ldots\cdot m_{n}(a_{n})}=\frac{1-m_{n}(a_{n})}{m_{n}(a_{n})},$
as desired. ∎
4. Examples of oscillating cocycles
In [TTD], Tserunyan and Tucker-Drob ask if it is possible to have a measure-class-preserving function $g:(Y,\nu)\rightarrow(Y,\nu)$ on a standard probability space $(Y,\nu)$ such that the Radon–Nikodym cocycle with respect to $\mu$, defined on the orbit equivalence relation induced by $g$, oscillates along the $g$-forward geodesic: that is $\limsup_{k\rightarrow\infty}\mathfrak{w}_{y}(g^{k}(y))=\infty$ and $\liminf_{k\rightarrow\infty}\mathfrak{w}_{y}(g^{k}(y))=0$ for a.e. $y\in Y$.
Letting $f:X\rightarrow X$ denote the least-deletion map defined above, we provide examples of product measures on $2^{\mathbb{N}}$ so that their Radon–Nikodym cocycles exhibit the desired oscillatory behaviour. Recall from above that $X$ is the subset of $2^{\mathbb{N}}$ of sequences with infinitely many 1s.
Example 4.1 (Oscillating cocycle).
Consider the product measure $\mu:=\prod_{n=0}^{\infty}m_{n}$ defined as follows. If $n\equiv 0$ mod $3$, set
$m_{n}(0):=1/3\text{ and }m_{n}(1):=2/3,$
so $m_{n}$ is biased towards $1$. If $n\not\equiv 0$ mod $3$, set
$m_{n}(0):=2/3\text{ and }m_{n}(1):=1/3,$
so $m_{n}$ is biased towards 00.
Theorem 4.2.
With the measure $\mu$ as above, the Radon–Nikodym cocycle $\mathfrak{w}$ oscillates along the forward geodesic of the least-deletion map $f$ for a.e. $x\in X.$
Proof.
Fix $x\in X$ and for each $k\geq 1,$ let $C_{k}(x):=\mathfrak{w}_{x}(f^{k}(x))$. By our computation of the Radon–Nikodym cocycle,
$C_{k}(x)=\prod_{i=1}^{k}\frac{m_{n_{i}}(0)}{m_{n_{i}}(1)}$
where $(n_{i})_{i\geq 1}$ is the subsequence of $\mathbb{N}$ for which $x_{n_{i}}=1$. Call $C(x):=(C_{k}(x))_{k\geq 0}$ the cocycle sequence at $x.$
We now define a "lazy" version of $C(x),$ denoted $\tilde{C}(x)$. We set $\tilde{C}_{0}(x):=1$, and define $\tilde{C}(x)$ inductively by
$\tilde{C}_{k+1}(x):=\begin{cases}\frac{m_{k}(0)}{m_{k}(1)}\tilde{C}_{k}(x)&\text{if }x_{k}=1\\ \tilde{C}_{k}(x)&\text{if }x_{k}=0\end{cases}$
Notice that $\tilde{C}_{k}$ jumps precisely when $k\in(n_{i})_{i\geq 1}$, and otherwise it lingers at the same value. We will prove that $\tilde{C}(x)$ oscillates and thus conclude that the cocycle sequence $C(x)$ oscillates.
For each $k\in\mathbb{N}$, observe that $\tilde{C}_{k}(x)$ is a power of 2. We will write $\tilde{C}_{k}$ instead of $\tilde{C}_{k}(x)$, since $x$ is fixed. Let $L_{k}:=\log_{2}\tilde{C}_{k}$. Then $L_{k}\in\mathbb{Z}$ for all $k$. Now set $X_{0}:=0$ and let
$Z_{k}:=L_{3k}-L_{3(k-1)}$
for $k>0$. Notice that $Z_{k}$ for $k\geq 1$ takes values in $\{-1,0,1,2\}$ deterministically according to the value of the sequence $x$ on the indices $3k-2,3k-1,$ and $3k$.
We now treat $x$ and $Z_{k}(x)$ as random variables, where $x$ is distributed according to $\mu$. Notice that
$Z_{k}(x)=\mathbbm{1}_{\{x_{3k-2}=1\}}+\mathbbm{1}_{\{x_{3k-1}=1\}}-\mathbbm{1}_{\{x_{3k}=1\}}.$
Since the product measure was periodically defined for triples, the $Z_{k}$ are independent identically distributed random variables.
By the linearity of expectation,
$\mathbb{E}[Z_{k}]$ $=\mathbb{P}(x_{3k-2}=1)+\mathbb{P}(x_{3k-1}=1)-\mathbb{P}(x_{3k}=1)=\frac{1}{3}+\frac{1}{3}-\frac{2}{3}=0.$
Since $Z_{k}=L_{3k}-L_{3(k-1)}$, we can write $L_{3k}=\sum_{i=1}^{k}Z_{i},$ so we have that $L_{3k}$ is a random walk on $\mathbb{Z}$ with zero bias. A theorem by Chung and Fuchs [chung] says that under mild conditions which are satisfied above, random walks on $\mathbb{Z}$ with zero bias oscillate almost surely.
Thus, for almost every sequence $x\in X,$
$\limsup_{k\to\infty}L_{3k}(x)=\infty\quad\text{and}\quad\liminf_{k\to\infty}L_{3k}(x)=-\infty,$
so exponentiating we find that $\mu$-a.e.,
$\limsup_{k\to\infty}\tilde{C}_{k}=\infty\quad\text{and}\quad\liminf_{k\to\infty}\tilde{C}_{k}=0.$
This gives us that the lazy cocycle sequence $\tilde{C}(x)$ oscillates almost surely. Thus $C(x)$ must also oscillate. ∎
Example 4.3 (An infinite family of oscillating cocycles).
We can generalize the above example to obtain a family of periodic measures, all of which have oscillating cocycles.
Fix $j>2$ and consider the infinite product measure $\mu:=\prod_{n=0}^{\infty}m_{n}$ defined as follows. If $n\equiv 0$ mod $j$, set
$m_{n}(0):=\frac{1}{j}\text{ and }m_{n}(1):=\frac{j-1}{j},$
and if $n\not\equiv 0$ mod $j$, set
$m_{n}(0):=\frac{j-1}{j}\text{ and }m_{n}(1):=\frac{1}{j}.$
Let $\mathfrak{w}$ be the cocycle of $\mu$ with respect to the least-deletion map $f:X\rightarrow X.$ The proof of example 4.2 can be straightforwardly generalized to give the following corollary.
Corollary 4.4.
With the measure $\mu$ as above, the Radon–Nikodym cocycle $\mathfrak{w}$ oscillates along the forward geodesic of the least-deletion map $f$ for a.e. $x\in X.$
5. Example of cocycle vanishing nonsummably along the forward geodesic
In [TTD], Tserunyan and Tucker-Drob ask if it is possible to have a measure-class-preserving function $g:(Y,\nu)\rightarrow(Y,\nu)$ on a standard probability space $(Y,\nu)$ such that the Radon–Nikodym cocycle with respect to $\nu$, defined on the orbit equivalence relation induced by $g$, vanishes nonsummably along the forward geodesic. That is,
$\lim_{k\rightarrow\infty}\mathfrak{w}_{y}(g^{k}(y))=0\quad\text{and}\quad\sum_{k\in\mathbb{N}}\mathfrak{w}_{y}(g^{k}(y))=\infty\quad\text{ for $\nu$-a.e. }y\in Y.$
We provide an example of a cocycle approaching zero nonsummably, on the space $X\subseteq 2^{\mathbb{N}}$ of sequences with infinitely many 1s, with the least-deletion map $f:X\rightarrow X$.